# Continual Learning for Encoder-only Language Models via a Discrete Key-Value Bottleneck

### v1

@@ Abstract
Continual learning remains challenging across various natural language understanding tasks. When models are updated with new training data, they risk catastrophic forgetting of prior knowledge. In the present work, we introduce a discrete key-value bottleneck for encoder-only language models, allowing for efficient continual learning by requiring only localized updates. Inspired by the success of a discrete key-value bottleneck in vision, we address new and NLP-specific challenges. We experiment with different bottleneck architectures to find the most suitable variants regarding language, and present a generic discrete key initialization technique for NLP that is task independent. We evaluate the discrete key-value bottleneck in four continual learning NLP scenarios and demonstrate that it alleviates catastrophic forgetting. We showcase that it offers competitive performance to other popular continual learning methods, with lower computational costs.
Continual Learning for Encoder-only Language Models
via a Discrete Key-Value Bottleneck
Andor Diera Ulm University andor.diera@uni-ulm.de                        Lukas Galke University of Southern Denmark galke@imada.sdu.dk                        Fabian Karl Ulm University fabian.karl@uni-ulm.de                        Ansgar Scherp Ulm University ansgar.scherp@uni-ulm.de
@@ 1 Introduction
Large language models are receiving increasing attention from the public due to their impressive zero-shot and few-shot abilities in a wide range of tasks Brown et al. (2020). Yet, for easier tasks where there is enough training data for supervised fine-tuning e.g. text classification, using smaller encoder-only language models is still preferable due to their often superior performance and lower computational requirements (Yuan et al., 2023; Yu et al., 2023; Qorib et al., 2024). Compared to large general-purpose models, fine-tuned networks lack general portability to new conditions and have limited generalization beyond their training distribution. For many target applications in natural language processing (NLP), training and test data can have a difference in the underlying distribution (Hupkes et al., 2022) and in the case of continual learning, the input distribution can change over time (Wang et al., 2024). To mitigate these challenges, different changes to model architectures and training regimens have been proposed Biesialska et al. (2020); Ke and Liu (2022); Wang et al. (2024). While many of these methods achieve strong improvements in continual learning setups, they often require task-specific modules and computationally demanding extensions to the base model (Ke et al., 2021; Buzzega et al., 2020).
In this work, we propose an adaptation of the Discrete Key-Value Bottleneck (DKVB) architecture Träuble et al. (2023) to the field of NLP. Discretization techniques can improve generalization in neural networks without introducing new task-specific parameters, regularization functions, or memory buffers (Liu et al., 2021, 2023; Träuble et al., 2023). More specifically, the DKVB architecture has shown strong performance in low-resource, class incremental learning scenarios in computer vision. This is due to local, context-dependent updates on learnable discrete key-value pairs that prevent catastrophic forgetting in the models.
To address the challenges of adapting DVKB to the field of NLP, we first analyze the effects of combining different variants of the DKVB with pre-trained encoder-only language models in standard learning scenarios. Subsequently, we take the best-performing DKVB configurations and evaluate their performance on continual learning scenarios. Finally, we show that given a dictionary of discrete keys optimized on a general-purpose corpus, DKVB offers competitive performance to state-of-the-art continual learning methods in most scenarios, but requires much lower training runtimes. The main contributions of our paper are:
  • We analyze different key optimization techniques and architectural variants of DKVB in the domain of NLP.
@@
  • We demonstrate that DKVB alleviates catasthropical forgetting and is more efficient compared to most other continual learning methods.
@@ 2.1 Continual Learning
Sequentially learning multiple tasks remains a significant challenge in the field of deep learning. Standard neural networks trained on a new task tend to forget most of the knowledge tied to tasks they have previously learned, leading to the phenomenon commonly labeled as catastrophic forgetting (McCloskey and Cohen, 1989; Van de Ven and Tolias, 2019). On the other hand, leveraging knowledge learned from old tasks to improve performance on new tasks, known as knowledge transfer, is a highly sought-after capability in NLP (Ke and Liu, 2022). Since re-training a model from scratch is often expensive, various methods for continual learning have been proposed to handle these challenges. Existing approaches in continual learning can be categorized into five distinct families: regularization-based, optimization-based, replay-based, architecture-based and instruction-based, with the latter being specific to large language models (Biesialska et al., 2020; Ke and Liu, 2022; Wang et al., 2024). A detailed description of these approaches can be found in Appendix A.1.
@@ 2.2 Discrete Representation Learning
Employing discrete variables in deep learning is challenging as indicated by the prevalence of continuous latent variables in most research papers, even when the underlying modality inherently involves discrete elements (e. g., text data). Van Den Oord et al. (2017) were the first to show the viability of large-scale discrete neural representation learning through the use of vector quantization. Their Vector Quantized-Variational Autoencoder (VQ-VAE) model utilizes a discrete latent space and thus avoids the “posterior collapse” problem common in many VAE models when the decoder ignores the latent space of the encoder and relies solely on the autoregressive properties of the input samples (Goyal et al., 2017). Subsequently, their methodologies have been widely employed in various applications, including audio (Borsos et al., 2023), videos (Yan et al., 2021), and anomaly detection (Marimont and Tarroni, 2021). More recently, discretization has been utilized for machine unlearning (Shah et al., 2023) and to improve disentangled representation learning (Noh et al., 2023) and robustness (Liu et al., 2021, 2023; Träuble et al., 2023). Discretization methods with bottlenecks have shown to improve generalization in reinforcement learning (Liu et al., 2021, 2023), visual reasoning (Liu et al., 2023), and vision-based continual learning (Träuble et al., 2023).
3 A Discrete Key-Value Bottleneck for Language Models
The DKVB architecture as described in Träuble et al. (2023) is fundamentally model and task-agnostic, but so far has been only studied in the field of computer vision. The use of a DKVB in language models poses new challenges that need to be addressed. These challenges include the sequential nature of the input data, the high dimensionality of the encoded representations, and the difference in commonly used pooling techniques between vision and language models. Below, we describe the base architecture of DKVB, alongside the process of key initialization and our proposed architectural adaptations and pre-experiments for finding the most suitable architectural variant.
@@ 3.1 Base Architecture and Key Initialization
The DKVB architecture (Träuble et al., 2023) follows a three-step process: encode, process via a discrete bottleneck, and decode. Figure 1 shows an overview of the base DKVB architecture.
@@
In the first step, an encoder model is used to project input vector $x$ into a lower dimensional vector $z\in\mathbb{R}^{m_{z}}$. This is followed by pooling (if needed) and partitioning $z$ into $C$ separate heads of dimension $d_{key}$. Each head possesses a unique discrete key-value codebook of size $K$, where the keys are initialized before training and are mapped to randomly initialized trainable value codes. In the second step, each head is first quantized by fetching the closest key (based on L2 distance) from the corresponding head’s codebook. Subsequently, the corresponding value code of dimension $d_{value}$ is retrieved for each head. Note, that the size of the bottleneck in respect to trainable parameters scales with the number of heads and codebook size. In the last step, the values are passed to a decoder to produce the final output. The decoder can be either parametric (with trainable weights) or non-parametric (by applying the softmax function to the mean pooled value codes).
The discrete keys of the bottleneck are initialized before training. Due to the 1-1 mapping between the keys and value codes, there is no gradient back-propagation between the values and keys. To ensure that the keys are broadly distributed in the feature space and have good representational power for given downstream tasks, they are first randomly initialized and then modified by using the encoded input samples as the basis for applying exponential moving average (EMA) updates (Van Den Oord et al., 2017). Alternatively, the keys can be initialized on input data different from the one in training, albeit with some decrease in downstream task performance (Träuble et al., 2023). After initialization, the keys are frozen and are not influenced by later changes in the input distribution shifts.
@@ 3.2 Architecture Adaptations for NLP
We introduce an adaptation of the DKVB architecture for the specific challenges in natural language processing. As argued above, these challenges are related to the high dimensionality of the data, pooling techniques, and decoding. To find the most fitting bottleneck architecture variants for the field of NLP, we conduct pre-experiments with different architectures. These architectural variants are motivated by the following NLP challenges.
@@
While natural language has an inherently discrete symbolic representation, text embeddings and language models using deep learning encode these discrete symbols into a continuous latent space (Muennighoff et al., 2023). This results in a high dimensional output $z\in\mathbb{R}^{t\times h}$, where $t$ is the token dimension (i.e., the number of tokens in the fixed length input sequence) and $h$ is the hidden dimension. Previous experiments with DKVB were conducted on low dimensional image data that has been pooled before forwarding output $z\in\mathbb{R}^{h}$ to the bottleneck (Träuble et al., 2023).
To address this difference, we design model variants with pooling applied before or after the bottleneck. Similarly, we experiment with creating the heads by partitioning hidden dimension $h$ and token dimension $t$ separately.
@@
Most modern convolutional networks in computer vision utilize max pooling in their pooling operations (He et al., 2015). Max pooling retains the most important features in image-related tasks but is less commonly used in NLP due to the loss of sequential information. The two most commonly used pooling techniques in NLP are mean pooling and pooling based on the CLS token. In mean pooling the contextualized token embeddings are averaged out, while the CLS pooling utilizes a special token optimized to represent the whole sequence (Devlin et al., 2018). We include both variants in our architecture search.
@@
Decoders with adjustable weights offer more expressiveness than non-parametric decoders but are more sensitive to changes in the training conditions (Ostapenko et al., 2022). For simple tasks where a linear mapping is sufficient, using just a softmax function as a non-parametric decoder might be appropriate. However, for many NLP tasks, it is crucial to capture complex patterns in the encoded representations Wang et al. (2018). We include both approaches in our experiments. For the parametric decoder, we concatenate the value codes and feed them into a simple linear layer preceded by a dropout layer. In the non-parametric version, we apply mean pooling on the values codes and apply a softmax function on the pooled representation.
@@ 3.3 Analyzing DKVB Variants for NLP
We analyze different variants of the DKVB architecture in encoder-only language models. For the pre-experiments, we use two popular text classification datasets. The R8 dataset which is a subset of the R21578 news dataset (Lewis, 1997) with 8 classes, and the Twenty Newsgroup (20ng) (Lang, 1995) which contains documents categorized into 20 newsgroups. For both datasets, we apply the standard train-test split, as used in Galke and Scherp (2022). We use a frozen BERT model as the pre-trained encoder for DKVB and perform a hyper-parameter search on the number of epochs, batch size, and learning rates. We report the performance of the best configurations. For learning rates, we found it is beneficial to have a high learning rate for the values layer. Additionally, in the case of the parametric decoder setup, a lower learning rate is applied to the decoder. For the discrete bottleneck parameters, we use a key dimension of $12$ and the number of key-value pairs of $4,096$ as in  Träuble et al. (2023). Key initialization is done before training for three epochs with an EMA decay of 0.2. Alongside the different variants for the DKVB, we list the results of a fine-tuned BERT and a frozen BERT with a fine-tuned linear classifier on top for reference. This we consider as the upper bounds. The test performance of the different architecture configurations can be seen in Table 1. The gap between the best-performing DKVB architecture and the fully fine-tuned BERT model is 2% on R8 and 7% on 20ng. The frozen BERT model achieved the same performance on R8 but attained 5% lower accuracy on 20ng compared to the best-performing DKVB variant. Overall the best performance was obtained by using the parametric decoder, applying mean pooling after the bottleneck, and using the hidden dimension as the base of the segmentation. To investigate the performance on other encoder-only language models, we experimented with RoBERTa Liu et al. (2019) and DistilBERT Sanh et al. (2019), and found the optimal bottleneck architecture to be the same (see Appendix A.4).
@@ 4 Continual Learning Settings
The goal of continual learning (CL) is to sequentially learn a function $f:X_{k}\rightarrow Y_{k}$ for all task $k$ in sequence $K$. Each task $k$ has a training set $M_{k}=\{(x_{i},y_{i},d_{i},t)\}^{N_{k}}_{i=1}$, where $x_{i}\in X_{k}$ is a training sample, $y_{i}\subseteq Y_{k}$ is a set of class labels, $d_{i}\subseteq D_{k}$ is the corresponding domain set (e. g., legal documents, movie reviews, news articles) of the sample, $t\in T_{k}$ is the task-type of the training set (e. g., sentiment analysis, topical classification, natural language inference etc), and $N_{k}$ is the number of samples in task $k$. To evaluate the DKVB architecture, we define three different incremental learning setups based on these components.
Domain Incremental Learning (DIL)
In the setting of DIL, the task type and class labels are assumed to be consistent across all tasks. The domain of the input changes between tasks, with each task having a set of non-overlapping domains $D_{k}\cap D_{k^{\prime}}=\emptyset$. A common DIL task-type in NLP is sentiment classification, where all tasks have the same class labels (i.e., positive, negative, neutral), but include samples from different source domains.
Class Incremental Learning (CIL)
In this setting, each task has a set of non-overlapping classes $Y_{k}\cap Y_{k^{\prime}}=\emptyset$. During testing, any previously learned class may be presented. CIL is generally considered to be the most challenging incremental learning scenario (Ke and Liu, 2022; Träuble et al., 2023). Apart from catastrophic forgetting and knowledge transfer, this setting includes the added complication of inter-task class separation (Kim et al., 2022). Inter-task class separation requires learning decision boundaries between the new task’s classes and the classes from previous tasks, all while lacking access to data from those previous tasks.
Task-type Incremental Learning (TIL)
The main challenge in this setting lies in the varying task-types. While it is possible that the tasks also have non-overlapping input domains and class labels in these settings, what differentiates TIL from other incremental learning scenarios are the disjoint task-types in each task $t_{k}\neq t_{k^{\prime}}$. This task-type is not identical to the type of objective function used in training (i. e., classification loss or regression loss), rather it defines the downstream task of the model, such as topical classification, sentiment analysis, or measuring semantic similarity. The scenario of using one model to learn different task-types has also been heavily researched in the field of multi-task learning (Crawshaw, 2020). The dominant approach in TIL is using a multi-head configuration with a separate head (or output layer) for each task. Since this decreases the probability of catastrophic forgetting, the main challenge in TIL is bi-directional knowledge transfer (Ke and Liu, 2022).

### v2

# Efficient Continual Learning for Small Language Models with a Discrete Key-Value Bottleneck

@@ Abstract
Continual learning remains a challenge across various natural language processing (NLP) tasks, as models updated with new training data often risk catastrophic forgetting of previously acquired knowledge. We introduce a discrete key-value bottleneck (DKVB) for encoder-only language models, enabling efficient continual learning through localized updates. Inspired by a discrete key-value bottleneck in vision, we consider new and NLP-specific challenges. We compare different bottleneck architectures for NLP and introduce a new, task-independent initialization technique for the discrete keys. We evaluate our DKVB for NLP in four continual learning scenarios and show that it alleviates catastrophic forgetting. Our experiments demonstrate that the proposed approach achieves competitive performance compared to popular continual learning methods while incurring lower computational costs. Furthermore, we show that DKVB remains effective even in challenging single-head continual learning scenarios where no task ID is provided. [^1] Source code available at: github.com/drndr/dkvb_nlp
@@ 1 Introduction
Large language models are receiving increasing attention from the public due to their impressive zero-shot and few-shot abilities in a wide range of tasks Brown et al. (2020). Yet, for easier tasks where there is enough training data for supervised fine-tuning, e. g., text classification, using smaller encoder-only language models is still preferable due to their often superior performance and lower computational requirements Yuan et al. (2023); Yu et al. (2023); Qorib et al. (2024); Li et al. (2025). Compared to large general-purpose models, fine-tuned networks lack general portability to new conditions and have limited generalization beyond their training distribution Luo et al. (2023). For many target applications in natural language processing (NLP), training and test data can have a difference in the underlying distribution Hupkes et al. (2023), and in the case of continual learning, the input distribution can change over time Wang et al. (2024). To mitigate these challenges, different changes to model architectures and training regimens have been proposed Biesialska et al. (2020); Ke and Liu (2022); Wang et al. (2024). While many of these methods improve continual learning, they often require task-specific modules and computationally demanding extensions to the base model Ke et al. (2021); Buzzega et al. (2020); Momeni et al. (2025).
In this work, we propose an adaptation of the Discrete Key-Value Bottleneck (DKVB) architecture Träuble et al. (2023) to the field of NLP. Discretization techniques can improve generalization in neural networks without introducing new task-specific parameters, regularization functions, or memory buffers Liu et al. (2021, 2023); Träuble et al. (2023). More specifically, the DKVB architecture has shown strong performance in low-resource, class incremental learning scenarios for computer vision. This is due to local, context-dependent updates on learnable discrete key-value pairs that prevent catastrophic forgetting in the models.
To address the challenges of adapting DKVB to NLP, we begin by analyzing how different variants of the discrete key-value bottleneck interact with pre-trained encoder-only language models in standard learning scenarios. In doing so, we tackle key challenges such as the high dimensionality of text representations, the choice of pooling strategies, and the design of an effective decoder head. Subsequently, we take the best-performing DKVB configurations and evaluate their performance in continual learning scenarios. Finally, we show that given a dictionary of discrete keys optimized on a general-purpose corpus, DKVB achieves similar effectiveness compared to leading continual learning approaches while requiring less training time. The main contributions of our paper are:
  • We analyze different optimization techniques and architectures of a DKVB in NLP using BERT, RoBERTa, and DistillBERT.
@@
  • We demonstrate that the DKVB alleviates catastrophic forgetting and is more efficient than most continual learning methods.
@@ 2.1 Continual Learning
Sequentially learning multiple tasks remains a significant challenge in the field of deep learning. Standard neural networks trained on a new task tend to forget most of the knowledge tied to tasks they have previously learned, leading to the phenomenon commonly labeled as catastrophic forgetting McCloskey and Cohen (1989); Van de Ven and Tolias (2019). On the other hand, leveraging knowledge learned from old tasks to improve performance on new tasks, known as knowledge transfer, is a highly sought-after capability in NLP Ke and Liu (2022). Since re-training a model from scratch is often expensive, various methods for continual learning have been proposed to handle these challenges. Existing approaches in continual learning can be categorized into five distinct families: regularization-based, optimization-based, replay-based, architecture-based, and instruction-based, with the latter being specific to large language models Biesialska et al. (2020); Ke and Liu (2022); Wang et al. (2024); Shi et al. (2024). A detailed description of these approaches can be found in Appendix A.
@@ 2.2 Discrete Representation Learning
Employing discrete variables in deep learning is challenging, as indicated by the prevalence of continuous latent variables in most research methods, even when the underlying modality inherently involves discrete elements (e. g., text data). Van Den Oord et al. (2017) were the first to show the viability of large-scale discrete neural representation learning through the use of vector quantization. Their Vector Quantized-Variational Autoencoder (VQ-VAE) model utilizes a discrete latent space and thus avoids the “posterior collapse” problem common in many VAE models when the decoder ignores the latent space of the encoder and relies solely on the autoregressive properties of the input samples Goyal et al. (2017). Subsequently, their methodologies have been widely employed in various applications, including audio Borsos et al. (2023), videos Yan et al. (2021), and anomaly detection Marimont and Tarroni (2021). More recently, discretization has been utilized for machine unlearning Shah et al. (2023) and to improve disentangled representation learning Noh et al. (2023) and robustness Liu et al. (2021, 2023); Träuble et al. (2023). Discretization methods with bottlenecks have been shown to improve generalization in reinforcement learning Liu et al. (2021, 2023), visual reasoning Liu et al. (2023), and vision-based continual learning Träuble et al. (2023).
3 A Discrete Key-Value Bottleneck for Encoder-only Language Models
The DKVB architecture as described in Träuble et al. (2023) is fundamentally model and task-agnostic, but so far has been only studied in the field of computer vision. The use of DKVB in language models poses new challenges, including the (i) sequential nature of the input data, the (ii) high dimensionality of the encoded representations, and the (iii) difference in commonly used pooling techniques between vision and language models. Below, we describe DKVB’s base architecture, the key initialization process, and our proposed architectural adaptations and pre-experiments for finding the most suitable architectural variant.
@@ 3.1 Base Architecture and Key Initialization
The DKVB architecture follows three steps: encode input, process via a discrete bottleneck, and decode. Figure 1 show an overview of the architecture.
@@
In the first step, an encoder model projects input vector $x$ into a lower dimensional vector $z\in\mathbb{R}^{m_{z}}$. This is followed by pooling (if needed) and partitioning $z$ into $C$ separate heads of dimension $d_{key}$. Each head possesses a unique discrete key-value codebook of size $K$, where the keys are initialized before training and are mapped to randomly initialized trainable value codes. In the second step, each head is first quantized by fetching the closest key (based on L2 distance) from the corresponding head’s codebook. Subsequently, the corresponding value code of dimension $d_{value}$ is retrieved for each head. Note the size of the bottleneck with respect to trainable parameter scales with the number of heads and codebook size. In the last step, the values are passed to a decoder to produce the final output. The decoder can be either parametric (with trainable weights) or non-parametric (by applying the softmax function to the mean pooled value codes).
The discrete keys of the bottleneck are initialized before training. Due to the 1-1 mapping between the keys and value codes, there is no gradient back-propagation between the values and keys. To ensure that the keys are broadly distributed in the feature space and have good representational power for given downstream tasks, they are first randomly initialized and then modified by using the encoded input samples as the basis for applying exponential moving average (EMA) updates Van Den Oord et al. (2017). Alternatively, the keys can be initialized on input data different from the one in training, albeit with some decrease in downstream task performance Träuble et al. (2023). After initialization, the keys are frozen and are not influenced by later changes in the input distribution shifts.
@@ 3.2 Architecture Adaptations for NLP
We introduce an adaptation of the DKVB architecture for the specific challenges in natural language processing. As argued above, these challenges are related to the high dimensionality of the data, pooling techniques, and decoding. We conduct pre-experiments with different architectures to find the most suitable bottleneck architecture variants and consider the following NLP-specific challenges:
@@
While natural language has an inherently discrete symbolic representation, text embeddings encode these discrete symbols into a continuous latent space Muennighoff et al. (2023). This results in a high dimensional output $z\in\mathbb{R}^{t\times h}$, where $t$ is the token dimension (i.e., the number of tokens in the fixed length input sequence) and $h$ is the hidden dimension. Previous experiments with DKVB were conducted on low dimensional image data that has been pooled before forwarding output $z\in\mathbb{R}^{h}$ to the bottleneck Träuble et al. (2023). To address this difference, we design model variants with pooling applied before or after the bottleneck. Similarly, we experiment with creating the heads by partitioning hidden dimension $h$ and token dimension $t$ separately.
@@
Most modern convolutional networks in computer vision utilize max pooling as pooling operation He et al. (2015). Max pooling retains the most important features in images but is less commonly used in NLP due to the loss of sequential information. The two most commonly used pooling techniques in NLP are mean pooling and pooling based on a special token (CLS). In mean pooling, the contextualized token embeddings are averaged out, while the CLS pooling utilizes a special token optimized to represent the whole sequence Devlin et al. (2018). We include both variants in our architecture search.
@@
Decoders with adjustable weights offer more expressiveness than non-parametric decoders but are more sensitive to changes in the training conditions Ostapenko et al. (2022). For simple tasks where linear mapping is sufficient, using just a softmax function as a non-parametric decoder might be appropriate. However, for many NLP tasks, it is crucial to capture complex patterns in the encoded representations Wang et al. (2018). We include both approaches in our experiments. For the parametric decoder, we concatenate the value codes and feed them into a simple linear layer preceded by a dropout layer. In the non-parametric version, we apply mean pooling on the values and apply a softmax function on the pooled representation.
@@ 3.3 Analyzing DKVB Variants for NLP
We analyze different variants of the DKVB architecture in encoder-only language models. For the pre-experiments, we use two popular text classification datasets. The R8 dataset, which is a subset of the R21578 news dataset Lewis (1997) with 8 classes, and the Twenty Newsgroup (20ng) Lang (1995) which contains documents categorized into 20 newsgroups. We apply the standard train-test split for both datasets, as used in Galke and Scherp (2022). We first use a frozen BERT model as the pre-trained encoder for DKVB and perform a hyperparameter search on the number of epochs, batch size, and learning rates.
We report the performance of the best configurations. For learning rates, we found it is beneficial to have a high learning rate for the values layer. Additionally, in the case of the parametric decoder setup, a lower learning rate is applied to the decoder. We use a key dimension of 121212 and the number of key-value pairs of 4,0964,0964 , 096 for the discrete bottleneck parameters as in  Träuble et al. (2023). Key initialization is done before training for three epochs with an EMA decay of 0.2. Alongside the different variants for the DKVB, we list the results of a fine-tuned BERT and a frozen BERT with a fine-tuned linear classifier on top for reference. This we consider as the upper bounds.
The test performance of the different architecture configurations can be seen in Table 1. The gap between the best-performing DKVB architecture and the fully fine-tuned BERT model is 2% on R8 and 7% on 20ng. The frozen BERT model achieved the same performance on R8 but attained 5% lower accuracy on 20ng compared to the best-performing DKVB variant. Overall the best performance was obtained by using the parametric decoder, applying mean pooling after the bottleneck, and using the hidden dimension as the base of the segmentation. To investigate the performance on other encoder-only language models, we experimented with RoBERTa Liu et al. (2019) and DistilBERT Sanh et al. (2019), and found the optimal bottleneck architecture to be the same (see Appendix C).
@@ 4 Continual Learning Settings
The goal of continual learning (CL) is to sequentially learn a function $f:X_{k}\rightarrow Y_{k}$ for all tasks $k$ in sequence $K$. Each task $k$ has a training set $M_{k}=\{(x_{i},y_{i},d_{i},t)\}^{N_{k}}_{i=1}$, where $x_{i}\in X_{k}$ is a training sample, $y_{i}\subseteq Y_{k}$ is a set of class labels, $d_{i}\subseteq D_{k}$ is the corresponding domain set (e. g., legal documents, movie reviews, news articles) of the sample, $t\in T_{k}$ is the task-type of the training set (e. g., sentiment analysis, topical classification, natural language inference etc), and $N_{k}$ is the number of samples in task $k$. To evaluate the DKVB architecture, we define three different incremental learning setups based on these components.
In the Domain Incremental Learning (DIL) setting, the task type and class labels are assumed to be consistent across all tasks. The domain of the input changes between tasks, with each task having a set of non-overlapping domains $D_{k}\cap D_{k^{\prime}}=\emptyset$. A common DIL task-type in NLP is sentiment classification, where all tasks have the same class labels (i. e., positive, negative, neutral), but include samples from different source domains.
In the Class Incremental Learning (CIL) setting, each task has a set of non-overlapping classes $Y_{k}\cap Y_{k^{\prime}}=\emptyset$. During testing, any previously learned class may be presented. CIL is generally considered the most challenging incremental learning scenario Ke and Liu (2022); Träuble et al. (2023). Apart from catastrophic forgetting and knowledge transfer, this setting includes the added complication of inter-task class separation Kim et al. (2022). Inter-task class separation requires learning decision boundaries between the new task’s classes and the classes from previous tasks without access to data from those previous tasks.
The main challenge in the Task-type Incremental Learning (TIL) setting lies in the varying task-types. While it is possible that the tasks also have non-overlapping input domains and class labels in these settings, what differentiates TIL from other incremental learning scenarios are the disjoint task-types in each task $t_{k}\neq t_{k^{\prime}}$. This task-type is not identical to the type of objective function used in training (i. e., classification loss or regression loss); rather, it defines the downstream task of the model, such as topical classification, sentiment analysis, or measuring semantic similarity. The scenario of using one model to learn different task-types has also been heavily researched in the field of multi-task learning Crawshaw (2020). The dominant approach in TIL is using a multi-head configuration with a separate head (or output layer) for each task. Since this decreases the probability of catastrophic forgetting, the main challenge in TIL is bi-directional knowledge transfer Ke and Liu (2022).